OpenStudy (anonymous):

Six subtracted from three times a number is -93. What would this equation look like?

OpenStudy (calculusxy):

(3*x)-6= -93

OpenStudy (anonymous):

Would X be -29?

OpenStudy (calculusxy):

Gimme a minute

OpenStudy (anonymous):

6-3x=-93

OpenStudy (calculusxy):

Why don't you do this. What is 6 subtracted from 93?

OpenStudy (anonymous):

-87

OpenStudy (calculusxy):

Okay then now divide 87/3.

OpenStudy (anonymous):

@savannaxx_ Wouldn't the 6 go after, not before?

OpenStudy (anonymous):

yes. very true.

OpenStudy (anonymous):

So it would be, 3x-6=-93

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

yeah, i think so

OpenStudy (calculusxy):

So for that i think that x holds the value of -29 becuase 3*(-29)=-87 and then (-87)-6=-93

OpenStudy (anonymous):

@calculusxy your making this way more complicated then it is

OpenStudy (anonymous):

Ok I think I got it, thanks to both of you!

OpenStudy (anonymous):

no problem :)

OpenStudy (calculusxy):

ok then people have different strategies so as long as you get the correct answer it will work @savannaxx_

OpenStudy (anonymous):

ok but why state the hardest strategy when the easy one will help people? im not being rude or anything im just wondering @calculsxy

OpenStudy (calculusxy):

Well thought that one of his strategy is going to work.